Which inverse operation would I use..? The equation is: 2x + 1/2 = 7/2
I tried subtracting 1/2 from both sides but I don't think my answer is right... Or would I divide 2 from2x and 7/2????

Answer by Alan3354(69443)

2x + 1/2 = 7/2
Subtract 1/2
2x = 3
Divide by 2
x = 3/2